# 3M-Diffusion: Latent Multi-Modal Diffusion for Text-Guided Generation of Molecular Graphs

### v1

@@ Abstract
Generating molecules with desired properties is a critical task with broad applications in drug discovery and materials design. Inspired by recent advances in large language models, there is a growing interest in using natural language descriptions of molecules to generate molecules with the desired properties. Most existing methods focus on generating molecules that precisely match the text description. However, practical applications call for methods that generate diverse, and ideally novel, molecules with the desired properties. We propose 3M-Diffusion, a novel multi-modal molecular graph generation method, to address this challenge. 3M-Diffusion first encodes molecular graphs into a graph latent space aligned with text descriptions. It then reconstructs the molecular structure and atomic attributes based on the given text descriptions using the molecule decoder. It then learns a probabilistic mapping from the text space to the latent molecular graph space using a diffusion model. The results of our extensive experiments on several datasets demonstrate that 3M-Diffusion can generate high-quality, novel and diverse molecular graphs that semantically match the textual description provided. Code is released at https://github.com/huaishengzhu/3MDiffusion.
Machine Learning, ICML
@@ 1 Introduction
Generating molecules with the desired properties is a critical task with broad applications in drug discovery and materials design (Hajduk & Greer, 2007; Mandal et al., 2009; Pyzer-Knapp et al., 2015). There is a growing interest in generative models to automate this task (You et al., 2018; Jin et al., 2018, 2020; Bjerrum & Threlfall, 2017).  Kusner et al. (2017) proposed a variational autoencoder that uses parse trees for probabilistic context free grammars to encode discrete sequences that specify the Simplified Molecular-Input Line-entry System (SMILES) (Weininger, 1988) notation of molecules use the resulting parse trees to generate SMILES codes that describe molecular graphs. Edwards et al. (2021) formulated molecule retrieval from textual description as a cross-lingual retrieval task. Recent advances in language models have led to innovative approaches to generate molecules from human-friendly textual descriptions of their properties, substructures, and biochemical activity (Edwards et al., 2022; Zeng et al., 2022; Zhao et al., 2023b; Fang et al., 2023a). These approaches learn mappings from textual descriptions to chemical structures using single transformer-based cross-lingual language model for molecule generation. Specifically, with the given text description of the molecule, these approaches produce SMILES strings describing the molecular graph.
Figure 1: The overview of 3M-diffusion, with a molecule graph encoder/decoder and a latent diffusion model conditioned on a prior (an aligned text encoder). The details of alignment in the text/graph encoders and diffusion model are given in Section 4.
@@
Despite promising initial results, such approaches suffer from some important limitations: (i) SMILES encoding is degenerate in that the a given molecular graph can have multiple SMILE encodings. For example, the structure of ethanol can be specified using CCO, OCC or C(O)C. Although a unique canonical SMILES representation can be produced for a given graph, the procedure used to do so relies on several arbitrary choices. Not surprisingly, molecules with identical chemical graphs may map to significantly distinct SMILES strings, making SMILES notation less than ideal for generating molecules from their textual descriptions (Jin et al., 2018). Important chemical properties corresponding to the substructures of molecules are more straightforward to express using molecular graphs rather than linear SMILES representations (Jin et al., 2018; Du et al., 2022); (ii) Using language models to generate diverse high-quality molecules matching a description is challenging, because the decoding process used such as beam search or greedy search often produces candidates that are often too similar to each other as shown in Figure 2. Hence, there is an urgent need for better approaches to generating diverse, novel molecular graphs from a given textual description  (Brown et al., 2019). Against this background, this paper aims to answer the following research question: How can we generate a diverse and novel collection of molecular graphs that match a given textual description?
@@
In this work, we introduce 3M-Diffusion, a novel multi-modal molecular diffusion method, for generating high-quality, diverse, novel molecular graphs given a textual description. Inspried by text-guided image generation using a diffusion model on the latent space (Rombach et al., 2022), 3M-Diffusion is trained to generate molecules from the latent distribution of a molecular graph autoencoder (See Figure 1). This enables the diffusion process to focus on the high-level semantics of molecules. We find that this approach is particularly well-suited for discrete molecular graph modality because it offloads the challenge of modeling a discrete and diverse distribution to the autoencoder and simplifies the diffusion process by restricting it to the continuous, latent feature space. Since graph data consists of both node attributes and graph structure, which differs significantly from the sequential and contextual nature of text data, to align the latent space of the text and molecular graph, 3M-diffusion first pretrains molecular graph and text encoders through contrastive learning on a large dataset of molecule-text pairs, enabling generate high-quality molecular graphs that match the given textual descriptions. Our extensive experiments demonstrate that 3M-Diffusion can generate high-quality, novel, and diverse molecular graphs that semantically match the textual description provided.
@@
Key Contributions. The key contributions of this paper are as follows: (i) To the best of our knowledge, 3M-Diffusion offers the first multimodal diffusion approach to generating molecular graphs from their textual descriptions, surpassing the limitations of state-of-the-art (SOTA) methods for this problem. (ii) 3M-Diffusion aligns the latent spaces of molecular graphs and textual descriptions to offer a text-molecule aligned latent diffusion model to generate higher-quality, diverse, and novel molecular graphs that match the textual description provided; and (iii) We present results of extensive experiments using four real-world text-based molecular graphs generation benchmarks that show that 3M-Diffusion outperforms SOTA methods for both text-guided and unconditional molecular graph generation. In particular, 3M-Diffusion achieves 146.27% novelty, 130.04% diversity relative improvement over the SOTA method with maintained semantics in textual prompt on the PCDes dataset.
@@ 2.1 Molecule Generation
Existing molecular graph generation methods can be broadly grouped into two categories: text-based and graph-based models. Text-based models typically utilize the SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2022) linear strings to describe each molecule  (Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018; Kusner et al., 2017; Flam-Shepherd et al., 2022; Fang et al., 2023b; Grisoni, 2023). Because a given molecular graph can have several linear string representations and slight changes to the string representation can yield significant changes in molecular graph being described, such linear string encodings are far from ideal for learning generative models for producing molecular graph from textual descriptions, several authors have used deep generative models, including graph variational autoencoders (VAEs) (You et al., 2018; Jin et al., 2018; Xiao et al., 2021; Jin et al., 2020; Kong et al., 2022; Liu et al., 2018; Diamant et al., 2023), normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021), generative adversarial networks (Guarino et al., 2017; Maziarka et al., 2020), diffusion models (Niu et al., 2020; Vignac et al., 2022; Jo et al., 2022) and autoregressive models (You et al., 2018; Xiao & Wang, 2021; Popova et al., 2019; Goyal et al., 2020) to learn distributions of molecular graphs from large databases of known graphs. However, many of these studies focus on generating molecular graphs with specified low-level properties, such as logP (the octanol-water partition coefficient) as opposed to those that match high-level textual descriptions of molecules (Edwards et al., 2022). However, practical applications, such as drug design, call for effective approaches to generating diverse and novel molecular graphs that match a given textual description.
2.2 Text-guided Generation of Molecular Graphs
@@
Advances in deep learning and language models have inspired a growing body of work on applications of such models to problems in the molecular sciences  (Schwaller et al., 2019; Xiao et al., 2023; Toniato et al., 2021; Xiao et al., 2024; Zhao et al., 2023a; Seidl et al., 2023; Schwaller et al., 2021; Vaucher et al., 2020). Advances in text-guided generation of images, videos, and audios  (Radford et al., 2021; Rombach et al., 2022; OpenAI, 2023; Yin et al., 2023), together with the flexibility offered by natural language descriptions of molecular graphs have inspired a growing body of work on text-guided generation of molecular graphs (Christofidellis et al., 2023; Edwards et al., 2022; Fang et al., 2023a; Zeng et al., 2022; Su et al., 2022). For example, MolT5 (Edwards et al., 2021, 2022) and ChemT5 (Christofidellis et al., 2023) pre-train on a substantial volume of unlabeled language text and SMILES strings for tasks such as molecule captioning and text-based molecule generation (Christofidellis et al., 2023; Edwards et al., 2022). However, as noted above, the degenerate nature of linear SMILES string representation of molecular graphs makes it less ideal for generating molecules from their textual descriptions. To address this problem, we propose 3M-Diffusion, aligning latent representations of textual descriptions for molecules with the latent representations of the graphs themselves, offering a powerful approach for text-guided molecule generation.
@@ 3.1 Problem Definition
In this paper, we consider generative modeling of 2D molecular graphs from scratch. Each molecule is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{|\mathcal{V}|}\right\}$ is the set of $|\mathcal{V}|$ nodes and $\mathcal{E}$ is the set of edges. Let $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{|\mathcal{V}\mid}\right]\in\mathbb{R}^{\left[\mathcal{V}\mid\times D_{x}\right.}$ be the node attribute matrix, where $x_{i}$ is the $D_{x}$-dimensional one-hot encoding feature vector of $v_{i}$, such as atomic type and chirality type. Similarly, a tensor $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times b}$ groups the one-hot encoding $\mathbf{e}_{ij}$ of each edge, with each entry representing a distinct edge type (bond types for molecule), while treating the absence of an edge as a specific edge type. Given the dataset $\mathcal{D}=\{(\mathcal{G}_{i},\mathcal{T}_{i})_{i=1}^{N}\}$ where molecule $\mathcal{G}$ has specific text descriptions $\mathcal{T}$ characterizing the features of the molecule, our goal is to learn conditional generation models $p_{\theta}(\mathcal{G}\mid\mathcal{T})$ for generating novel, diverse, and valid molecules while simultaneously aligning with the desired features outlined in text descriptions $\mathcal{T}$. For brevity, in what follows, we omit the input subscript $i$ unless otherwise stated.
@@ 3.2 Diffusion Models
$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{x}_{t},t)$ (2)
$+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$
@@ 4 Latent Multi-Modal Diffusion for Molecules
Our 3M-Diffusion in Figure 1 aims to learn a probabilistic mapping from the latent space of textual descriptions to a latent space of molecular graphs for text-guided generation of molecular graphs. However, direct mapping between these two latent representations using a diffusion model faces a key hurdle, namely, large mismatch between the latent spaces of textual descriptions and of molecular graphs. To overcome this hurdle, 3M-Diffusion adopts a two-part approach using a text-molecule aligned variational autoencoder and a multi-modal molecule latent diffusion. The former learns to align the representation of molecular graphs with that of their textual descriptions using contrastive learning, followed by a decoder that maps the latent representation back to the corresponding molecular graphs. The latter works with the text-aligned latent representation of molecular graphs to learn a conditional generative model, that maps textual descriptions to latent representations of molecular graphs. As we will see later, 3M-Diffusion is able to produce a diverse collection of novel molecular graphs that match the given textual descriptions.
@@ 4.1 Text-Molecule Aligned Variational Autoencoder
Our text-molecule aligned variational autonecoder comprises of three parts: the molecular graph encoder $E_{g}$, text encoder $E_{t}$ and molecular graph decoder $D$. To bridge the representation gap between molecular graphs and their textual descriptions, the text and molecular graph encoders are trained on a large-scale dataset of molecule-text pairs. These encoders learn a well-aligned shared text-molecule representation, enriching the semantics of the molecular graph representation by aligning it the latent representation of textual descriptins of molecules through contrastive learning. These three parts are described in more detail below:
@@
where $\bm{\mu}_{g}$ and $\bm{\sigma}_{g}$ represent the mean and standard deviations, respectively. Then, the latent graph representation $\mathbf{z}$ is sampled from a Gaussian distribution as $\mathbf{z}\sim q\left(\mathbf{z}\mid\mathbf{A},\mathbf{X}\right)\simeq\mathcal{N}\left(\bm{\mu}_{g},\bm{\sigma}_{g}\right)$. This graph encoding process can also be expressed as $\mathbf{z}=E_{g}(\mathcal{G})$ signifying the outcome of the encoding operation within the input context of $\mathcal{G}$.
Text Encoder maps the textual descriptions of molecular graphs into a latent space, that provides input to a conditional generative model of molecular graphs. This encoded information becomes instrumental in generating molecules that match a given textual description. To inject potentially useful scientific knowledge from the literature into $E_{t}$ we initialize it with Sci-BERT (Beltagy et al., 2019), an encoder-only transformer pretrained on scientific publications.
Aligning the representation of molecular graphs with that of their textual descriptions is crucial for establishing a shared representation for effective text-guided molecular graph generation. Graph data consists of both node attributes and graph structure, which differs significantly from the sequential and contextual nature of text data. We aim to align the molecule representation from the molecule graph encoder $E_{g}$ with text representation from the text encoder $E_{t}$ by pretraining. Consider a molecular graph $\mathcal{G}$ and its textual description $\mathcal{T}$. To align these two encoders and ensure a cohesive latent space, we employ the following contrastive learning loss over a batch $\mathcal{B}\in\mathcal{D}$ of large-scale dataset containing molecule-text pairs to pretrain encoders:
@@
where $\mathbf{z}_{i}=E_{g}(\mathcal{G}_{i})$, $\mathbf{c}_{i}=E_{t}(\mathcal{T}_{i})$ and $\mathbf{c}_{j}=E_{t}(\mathcal{T}_{j})$ are representations of corresponding molecule graph and texts, respectively. $\cos(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is the temperature hyperparameter. The encoders are pretrained on molecule-text (300K) pairs from PubChem collected by  Liu et al. (2023) and robust enough to capture a well-aligned molecule-text space, which overcomes the challenges posed by the distribution gap and produce more consistent and high-quality results in cross-modal molecule generation.
Molecular Graph Decoder learns to use latent representation $\mathbf{z}$ to decode the molecular graph. In this work, we employ the HierVAE (Jin et al., 2020) decoder, as it works empirically well in our final model. Note that our approach remains flexible and can be easily applied to other graph decoders (Li et al., 2018; Kong et al., 2022; Grover et al., 2019). Specifically, it generates a molecule $\mathcal{G}$ using a graph representation $\mathbf{z}$ from the encoder. Given the molecular graph $\mathcal{G}$, we minimize the negative Evidence Lower Bound (ELBO) (Jin et al., 2020) to train both encoder and decoder:
@@
Here $\hat{\mathcal{G}}=D(\mathbf{z})$, in which $D$ denotes the decoder parameterized by HierVAE (Jin et al., 2020). $\mathrm{KL}[q(\cdot)||p(\cdot)]$ is the Kullback-Leibler (KL) divergence between $q(\textperiodcentered\cdot)$ and $p(\cdot)$. We follow Jin et al. (2020) and set the weight of KL loss $\alpha=0.1$. Here, $p(\mathbf{z})=\mathcal{N}\left(\mathbf{z}\mid\mathbf{0},\mathbf{I}\right)$ is the isotropic multivariate Gaussian prior, where $\mathbf{I}$ is the identity matrix. We optimize this first expectation term concurrently through reparameterization trick (Kingma & Welling, 2013).

### v2

# 3M-Diffusion: Latent Multi-Modal Diffusion for Language-Guided Molecular Structure Generation

@@ Abstract
Generating molecular structures with desired properties is a critical task with broad applications in drug discovery and materials design. We propose 3M-Diffusion, a novel multi-modal molecular graph generation method, to generate diverse, ideally novel molecular structures with desired properties. 3M-Diffusion encodes molecular graphs into a graph latent space which it then aligns with the text space learned by encoder-based LLMs from textual descriptions. It then reconstructs the molecular structure and atomic attributes based on the given text descriptions using the molecule decoder. It then learns a probabilistic mapping from the text space to the latent molecular graph space using a diffusion model. The results of our extensive experiments on several datasets demonstrate that 3M-Diffusion can generate high-quality, novel and diverse molecular graphs that semantically match the textual description provided. The code is available on github.
**footnotetext: Equal contribution
@@ 1 Introduction
Generating molecular structures with the desired properties is a critical task with broad applications in drug discovery and materials design (Hajduk & Greer, 2007; Mandal et al., 2009; Pyzer-Knapp et al., 2015). There is a growing interest in generative models to automate this task (You et al., 2018; Jin et al., 2018; 2020; Bjerrum & Threlfall, 2017).  Kusner et al. (2017) proposed a variational autoencoder that uses parse trees for probabilistic context free grammars to encode discrete sequences that specify the Simplified Molecular-Input Line-entry System (SMILES) (Weininger, 1988) notation of molecules use the resulting parse trees to generate SMILES codes that describe molecular graphs. Edwards et al. (2021) formulated molecule retrieval from textual description as a cross-lingual retrieval task.
Recent advances in language models have led to innovative approaches to generate molecular structures that match human-friendly textual descriptions of their properties, substructures, and biochemical activity (Edwards et al., 2022; Zeng et al., 2022; Zhao et al., 2023b; Fang et al., 2023a). These approaches learn mappings from textual descriptions to molecular structures using single transformer-based cross-lingual language model for molecule generation. Specifically, with the given text description of a molecule, these approaches produce SMILES strings that encode the molecular structure.
@@
Despite promising initial results, such approaches suffer from some important limitations: (i) SMILES encoding is degenerate in that the a given molecular structure can have multiple SMILE encodings. For example, the structure of ethanol can be specified using CCO, OCC or C(O)C. Although in principle a unique canonical SMILES representation can be produced for a given molecular structure, the procedure used to do so relies on several arbitrary choices. Not surprisingly, molecules with identical structure may map to different SMILES strings. This makes SMILES notation less than ideal for generating molecules from their textual descriptions (Jin et al., 2018). Important chemical properties corresponding to the substructures of molecules are more straightforward to express using molecular graphs rather than linear SMILES representations (Jin et al., 2018; Du et al., 2022); (ii) Using language models to generate diverse and high-quality molecular structures matching a description is challenging, because the process used to decode the learned representation into the corresponding molecular structure, e.g., beam search or greedy search, often produces structures that are often too similar to each other (See Figure 2). Hence, there is an urgent need for better approaches for generating diverse, novel molecular structures from a given textual description  (Brown et al., 2019). Against this background, this paper aims to answer the following research question: How can we generate a diverse and novel collection of molecular graphs that match a given textual description?
Figure 1: The overview of 3M-diffusion, with a molecular graph encoder/decoder and a latent diffusion model conditioned on a prior (an aligned LLM encoder). The details of alignment in the text/graph encoders and diffusion model are given in Section 4.
@@
We introduce 3M-Diffusion, a novel multi-modal molecular diffusion method, for generating high-quality, diverse, novel molecular structures from a textual description. Inspired by text-guided image generation using a diffusion model on the latent space (Rombach et al., 2022), 3M-Diffusion is trained to generate molecular structures from the latent distribution of a molecular graph autoencoder (See Figure 1). This enables the diffusion process to focus on the high-level semantics of encodings of molecular structures. We find that this approach is particularly well-suited for discrete molecular graph modality because it relegates the challenge of modeling a discrete and diverse distribution to the autoencoder and simplifies the diffusion process by restricting it to the continuous, latent feature space. Since graph data consists of both node attributes and graph structure, which differs significantly from the sequential text data, to align the latent representations of textual descriptions and molecular graphs, 3M-diffusion employs contrastive learning on a large data set of pairs of molecular structures and their textual descriptions to pretrain molecular graph encoders and LLM text encoders. This enables 3M-diffusion to generate high-quality molecular graphs that match the given textual descriptions.
@@
Key Contributions. The key contributions of this paper are as follows: (i) To the best of our knowledge, 3M-Diffusion offers the first multimodal diffusion approach to generating molecular structures from their textual descriptions, surpassing the limitations of state-of-the-art (SOTA) methods for this problem. (ii) 3M-Diffusion aligns the latent spaces of molecular graphs and textual descriptions to offer a text-molecule aligned latent diffusion model to generate higher-quality, diverse, and novel molecular structures that match the textual description provided; and (iii) Results of extensive experiments using four real-world text-based molecular graphs generation benchmarks show that 3M-Diffusion outperforms SOTA methods for both text-guided and unconditional molecular structure generation. In particular, 3M-Diffusion achieves 146.27% novelty, 130.04% diversity relative improvement over the SOTA method with maintained semantics in textual prompt on the PCDes dataset.
Figure 2: Qualitative comparisons to the MolT5-large of generated molecules on ChEBI-20. Compared with the SOTA method MolT5-large, our results are more diverse and novel with maintained semantics in textual prompt. More results are provided in Appendix D.4
@@ 2.1 Molecular Structure Generation
Existing molecular graph generation methods can be broadly grouped into two categories: text-based and graph-based models. Text-based models typically utilize the SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2022) linear strings to describe each molecule  (Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018; Kusner et al., 2017; Flam-Shepherd et al., 2022; Fang et al., 2023b; Grisoni, 2023). Because a given molecular graph can have several linear string representations and small changes to the string representation can result in large changes in molecular graph being described, such linear string encodings are far from ideal for learning generative models for producing molecular graph from textual descriptions. Inspired by the success of learning enhanced graph representations using graph based models (Kipf & Welling, 2016; You et al., 2020; Xiao et al., ; Xiao et al., 2024), several authors have employed deep generative models, including graph variational autoencoders (VAEs) (You et al., 2018; Jin et al., 2018; Xiao et al., 2021; Jin et al., 2020; Kong et al., 2022; Liu et al., 2018; Diamant et al., 2023), normalizing flows (Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021), generative adversarial networks (Guarino et al., 2017; Maziarka et al., 2020), diffusion models (Niu et al., 2020; Vignac et al., 2022; Jo et al., 2022) and autoregressive models (You et al., 2018; Xiao & Wang, 2021; Popova et al., 2019; Goyal et al., 2020) to learn distributions of molecular graphs from large databases of known graphs. However, many of these studies focus on generating molecular graphs with specific low-level properties, such as logP (the octanol-water partition coefficient) as opposed to those that match high-level textual descriptions of a broader range of molecular properties  (Edwards et al., 2022), e.g., water solubility, chemical activity, etc. However, practical applications, such as drug design, call for effective approaches to generating diverse and novel molecular graphs that match a given textual description of such high-level properties.
2.2 Language-guided Molecule Generation
@@
Advances in deep learning and language models have inspired a growing body of work on applications of such models to problems in the molecular sciences  (Schwaller et al., 2019; Xiao et al., 2023; Toniato et al., 2021; Xiao et al., 2024; Zhao et al., 2023a; Seidl et al., 2023; Schwaller et al., 2021; Vaucher et al., 2020). Advances in text-guided generation of images, videos, and audios  (Radford et al., 2021; Rombach et al., 2022; OpenAI, 2023; Yin et al., 2023), together with the flexibility offered by natural language descriptions of molecular graphs have inspired a growing body of work on text-guided generation of molecular graphs (Christofidellis et al., 2023; Edwards et al., 2022; Fang et al., 2023a; Zeng et al., 2022; Su et al., 2022). For example, MolT5 (Edwards et al., 2021; 2022) and ChemT5 (Christofidellis et al., 2023) pre-train on a substantial volume of unlabeled language text and SMILES strings for tasks such as molecule captioning and text-based molecule generation (Christofidellis et al., 2023; Edwards et al., 2022). However, as noted above, the degenerate nature of linear SMILES string representation of molecular graphs makes it less ideal for generating molecules from their textual descriptions. To address this problem, we propose 3M-Diffusion, a novel method that aligns the latent representation of textual descriptions of molecules with that of the corresponding molecular structures, yielding a powerful approach to text-guided molecular structure generation.
@@ 3.1 Problem Definition
We consider generative modeling of 2D molecular graphs. Each molecule is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{|\mathcal{V}|}\right\}$ is the set of $|\mathcal{V}|$ nodes and $\mathcal{E}$ is the set of edges. Let $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{|\mathcal{V}\mid}\right]\in\mathbb{R}^{\left[\mathcal{V}\mid\times D_{x}\right.}$ be the node attribute matrix, where $x_{i}$ is the $D_{x}$-dimensional one-hot encoding feature vector of $v_{i}$, such as atomic type and chirality type. Similarly, a tensor $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times b}$ groups the one-hot encoding $\mathbf{e}_{ij}$ of each edge, where each entry is a distinct edge type (bond types for molecule), with the absence of an edge encoded explicitly using a designated edge type. Given the dataset $\mathcal{D}=\{(\mathcal{G}_{i},\mathcal{T}_{i})_{i=1}^{N}\}$ where molecule $\mathcal{G}$ has specific text descriptions $\mathcal{T}$ characterizing the features of the molecule, our goal is to learn conditional generation models $p_{\theta}(\mathcal{G}\mid\mathcal{T})$ for generating novel, diverse, and valid molecules while also aligning with the desired features outlined in text descriptions $\mathcal{T}$. For brevity, in what follows, we omit the input subscript $i$ when it is clear from context.
@@ 3.2 Diffusion Models
$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{x}_{t},t)+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$ (2)
@@ 4 Latent Multi-Modal Diffusion for Molecules
3M-Diffusion (See Figure 1) aims to learn a probabilistic mapping from the latent space of textual descriptions to a latent space of molecular graphs for text-guided generation of molecular graphs. However, direct mapping between these two latent representations using a diffusion model faces a key hurdle, namely, large mismatch between the latent spaces of textual descriptions and of molecular graphs. To overcome this hurdle, 3M-Diffusion adopts a two-step approach. The first step uses contrastive learning to produce text-molecule aligned variational autoencoder that align the representation of molecular graphs with that of their textual descriptions, and a decoder that maps the latent representation back to the corresponding molecular graphs. The second step uses the text-aligned latent representation of molecular graphs to learn a conditional generative model, that maps textual descriptions to latent representations of molecular graphs. As we will see later, 3M-Diffusion can produce diverse novel molecular graphs that match the given textual descriptions.
@@ 4.1 Text-Molecule Aligned Variational Autoencoder
The proposed text-molecule aligned variational autonecoder comprises of three parts: the molecular graph encoder $E_{g}$, LLM encoder $E_{t}$ and molecular graph decoder $D$. To bridge the representation gap between molecular graphs and their textual descriptions, the text and molecular graph encoders are trained on a large-scale dataset of molecule-text pairs. These encoders use contrastive learning to construct well-aligned text-molecular structure encodings that align the latent representation of molecular structures by aligning it with with that of their textual descriptions. The details of this process are given below.
@@
where $\bm{\mu}_{g}$ and $\bm{\sigma}_{g}$ represent the mean and standard deviations, respectively. Then, the latent graph representation $\mathbf{z}$ is sampled from a Gaussian distribution as $\mathbf{z}\sim q\left(\mathbf{z}\mid\mathbf{A},\mathbf{X}\right)\simeq\mathcal{N}\left(\bm{\mu}_{g},\bm{\sigma}_{g}\right)$. This graph encoding process can be expressed as $\mathbf{z}=E_{g}(\mathcal{G})$ signifying the outcome of the encoding operation in the input context of $\mathcal{G}$.
LLM Encoder maps the textual descriptions of molecular structures into a latent space that provides input to a conditional generative model of molecular graphs. This encoded information is instrumental in generating molecules that match a given textual description. To inject potentially useful scientific knowledge from the literature into $E_{t}$ we initialize it with encoder-based LLM, i.e., Sci-BERT (Beltagy et al., 2019), which is pretrained on the text of scientific publications.
Aligning the representation of molecular graphs with that of their textual descriptions is crucial for establishing a shared representation for effective text-guided molecular graph generation. Graph data encodes both node attributes and graph structure which differs significantly from the sequential and contextual nature of text data. We aim to align the molecular representation from the molecule graph encoder $E_{g}$ with text representation from the LLM encoder $E_{t}$ using contrastive learning. Consider a molecular graph $\mathcal{G}$ and its textual description $\mathcal{T}$. To align these two encoders and ensure a cohesive latent space, we employ the following contrastive learning loss over a batch $\mathcal{B}\in\mathcal{D}$ of large-scale dataset containing molecule-text pairs to pretrain encoders:
@@
where $\mathbf{z}_{i}=E_{g}(\mathcal{G}_{i})$, $\mathbf{c}_{i}=E_{t}(\mathcal{T}_{i})$ and $\mathbf{c}_{j}=E_{t}(\mathcal{T}_{j})$ are representations of corresponding molecule graph and texts, respectively. $\cos(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is the temperature hyperparameter. The encoders are pretrained on molecule-text (300K) pairs from PubChem  Liu et al. (2023). As we will see, this process yields a well-aligned molecule-text space that can support the generation of diverse and novel molecular structures from their textual descriptions.
Molecular Graph Decoder learns to use latent representation $\mathbf{z}$ to generate the corresponding molecular structure. In this work, we employ the HierVAE (Jin et al., 2020) decoder. Note however that our approach can be combined with other graph decoders (Li et al., 2018; Kong et al., 2022; Grover et al., 2019). Specifically, the molecular graph decoder generates a molecular structure $\mathcal{G}$ using a graph representation $\mathbf{z}$ from the encoder. Given the molecular graph $\mathcal{G}$, we minimize the negative Evidence Lower Bound (ELBO) (Jin et al., 2020) to train both the encoder and the decoder:
@@
Here $\hat{\mathcal{G}}=D(\mathbf{z})$, in which $D$ denotes the decoder parameterized by HierVAE (Jin et al., 2020). $\mathrm{KL}[q(\cdot)||p(\cdot)]$ is the Kullback-Leibler (KL) divergence between $q(\textperiodcentered\cdot)$ and $p(\cdot)$. We follow Jin et al. (2020) and set the weight of KL loss $\alpha=0.1$. Here, $p(\mathbf{z})=\mathcal{N}\left(\mathbf{z}\mid\mathbf{0},\mathbf{I}\right)$ is the isotropic multivariate Gaussian prior, where $\mathbf{I}$ is the identity matrix. We optimize this first expectation term concurrently using the reparameterization trick (Kingma & Welling, 2013).